OpenStudy (anonymous):

GRADE 10 QUESTION SYSTEM OF EQUATIONS DON'T SOLVE, JUST HELP ME PLEASE. 7. Tyler rows 10km downstream in 2 h. On the return trip, it takes him 4h to travel 8km. Determine his average rowing speed and the speed of the current.

OpenStudy (the_fizicx99):

So um 10km in 2 hours so per hour its 5km/h 8km in 4 hours, means 2km in 1 hour or per hour x + y = 5km and x - y = 2km Now you have a system, the Y's cancel out and you have 2x = 7km solve for x by dividing both sides by 2 2x = 7 /2 /2 x = 3.5 Now solve for y, 3.5 + y = 5 -3.5 -3.5 y = 1.5 So the average speed it 1.5km

OpenStudy (anonymous):

oh. sorry i think i closed it as soon as you replied by accident

OpenStudy (the_fizicx99):

>.<

OpenStudy (anonymous):

i still don tunderstanf

OpenStudy (anonymous):

isn't it x+y=18

OpenStudy (the_fizicx99):

Why 18? You're not adding lol >_<